Margin of safety as a percentage of sales

1.) Compute the company's margin of safety as a percentage of its sales??

Molander Coporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month's budget appear below:

Selling price = $30 per unit
Variable expenses = $20 per unit
Fixed expenses = $7,500 per month
Unit Sales = 1,000 units per month
Answer #1
Margin of Safety (in pesos) = 1000 units x $30 – 750 units x $30 = $7500
Margin of Safety (in units) = 1000 units – 750 units = 250 units

Selling Price $30
Variable Cost $20
Contribution Margin $10

Contribution Margin Ratio = $10 / $30 = 33.33%

Break – Even Sales (in pesos) = $7500 / 33.33% = $22500
Break – Even Sales (in units) = $22500 / $30 = 750 units

Margin of Safety (in pesos) = 1000 units x $30 – 750 units x $30 = $7500
Margin of Safety (in units) = 1000 units – 750 units = 250 units
answered by: Sheba
Answer #2
[Margin of Safety in dollars = Total budgeted or actual sales - Break even sales]

Actual Sales = ($30 per unit)(1,000 units per month) = $30,000

Break even sales = $7,500 + ($20 per unit)(1,000 units per month) = $27,500

Margin of Safety in dollars = $30,000 - $27,500 = $2500

Margin of Safety = $2500 / $30,000 = 8.33%
answered by: vanny :)